Grade 11Physical Chemistry

What are the necessary conditions for a compound to be aromatic ?

Tamanna
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
The necessary conditions for any system to be aromatic are:
(i) The molecule should be planar
(ii) It should be cyclic with alternate single and
double bonds and the cyclic $\pi $ cloud should surround all the carbon atoms of the ring.
(iii) It should contain (4n +2)$\pi $ electrons where n = 0,1,2,3, etc.
A molecule which does not satisfy any one or more of the above conditions is said to be non-aromatic.
 
 
Regards
Arun (askIITians forum expert)